OpenStudy (anonymous):

Please help, I am so confused! I need to find the simplest form of the expression: sqrt 2 - sqrt 10 / sqrt 2 + sqrt 10

OpenStudy (mathmale):

\[\sqrt 2 - \sqrt 10 / \sqrt 2 + \sqrt 10=\sqrt 2 - \sqrt (10/2) + \sqrt 10\] Then, sqrt 2 - sqrt (10/2) + sqrt 10=sqrt 2 - sqrt 5 + sqrt 10

OpenStudy (mathmale):

I'd regard this to be an adequate answer. Can't simplify it any further. Questions?

OpenStudy (anonymous):

So that equals 2.3404? I need to come up with one of the following answers: A.-3/2+1/4 sqrt10 B.-3/2+1/2sqrt5 C. 1-1/3sqrt5 D. 1+1/2sqrt5

OpenStudy (anonymous):

Not sure what I am doing wrong here...

OpenStudy (mathmale):

Please stick with those radicals. If you evaluate them with your calculator, none are going to match your four answer options.

OpenStudy (mathmale):

Same here: I'm not sure what I'm overlooking! You might want to try reviwing the original question so you can be certain you copied it correctly. You could also post this as a separate, new question: How might the following expression be simplified further? Sqrt(2) -Sqrt(5) + Sqrt(10). Wish i could help more, but I cannot do that at this point.

OpenStudy (anonymous):

ok thanks for trying (=